Fc=mv/r
Fc is centripetal force, m is mass, v is speed and r is radius


6. What will happen to the size of the centripetal force if the mass (m) is doubled?


1)No change
2)The centripetal force doubles
3)The centripetal force increases by a factor of 4
4)The centripetal force is halved
1
Expert's answer
2019-10-10T09:45:01-0400

As follows from the formula for centripetal force

"F_c=(mv^2)/r=\frac{v^2}{r}\cdot m"

centripetal force is directly proportional to mass

"F_c\backsim m"

Answer: 2)The centripetal force doubles